OpenStudy (anonymous):

solve 5e^(-x) - 7 = 2

OpenStudy (tkhunny):

Have you considered adding 7 to both sides?

OpenStudy (anonymous):

e^-x = 9/5 -x ln(e) = ln(9/5) x = ln(5/9) since ln(e) = 1

OpenStudy (tkhunny):

Gaaa!!!! Now, we'll never know if idkatie can do it!

OpenStudy (anonymous):

dude i just need to know the answer

OpenStudy (tkhunny):

That is NEVER correct. You need to LEARN! Stop underestimating yourself. Also, please read the forum rules, both of you.